OpenStudy (anonymous):

Geometry help, please! Will give medal 2) Write an equation of a line that passes through the point (8,4) and is parallel to the like y=4x+2 * y= 4x -28 * y= 4x + 28 * y= 1/4x -2 * y= 1/4 x +2

OpenStudy (anonymous):

Use graphing online by doing a Google search on Demoso Graphing

OpenStudy (anonymous):

Then graph each equation above and see if it matches what the question is asking

OpenStudy (anonymous):

@Jayeeeem

OpenStudy (anonymous):

since required line is parallel to y=4x+2 Comparing y=4x+2 by y= mx+c we find slope m = 4 So, Slope of the required line which is paralle to y=4x+2 would be = -1/4 Since line passes through (8,4). therefore eq of line in slope point form is given as; \[\huge y-y_1 = m(x-x_1)\] where x1= 8 and y1= 4 \[\huge \rightarrow y-4 =-\frac {1}{4}(x-8)\] \[\huge \rightarrow y-4 =-\frac {1}{4}x +2\] \[\huge \rightarrow y-4 =-\frac {1}{4}x +2 + 4\] \[\huge \rightarrow y =-\frac {1}{4}x +6\]

OpenStudy (anonymous):

Which is the required answer...

OpenStudy (anonymous):

@Jayeeeem

OpenStudy (amistre64):

hmm, parallel lines dont have perpendicular slopes

OpenStudy (amistre64):

the 'required' answer relies in the askers participation in finding the solution. otherwise, fancy writing of equations that are incorrectly applied to a problem simply doesnt help out.